OpenStudy (anonymous):

If x+12=6cos(theta) for 0

OpenStudy (anonymous):

since \(\large x+12=6cos\theta \rightarrow \frac{x+12}{6}=cos\theta \) . now, the double angle formula states: \(\large cos(2\theta)=cos^2\theta-sin^2\theta \). \(\large cos^2\theta=\frac{(x+12)^2}{6^2} \) \(\large sin^2\theta=1-cos^2\theta = 1-\frac{x+12)^2}{6^2}\) can you take it from here?

OpenStudy (anonymous):

ok ill try thanks much

OpenStudy (anonymous):

yw... lemme know if u get stuck.... :) btw.... that last line should read \(\large sin^2\theta=1-cos^2\theta=1-\frac{(x+12)^2}{6^2}\)

OpenStudy (anonymous):

i tried it but my answer in the webwork and it turn out to be wrong........well i actually used the same answer u gave me

OpenStudy (anonymous):

this is what i got.... (i skipped some algebra but it's just basically simplification): \(\large cos2\theta = cos^2\theta - sin^2\theta=\frac{(x+12)^2}{36}-(1-\frac{(x+12)^2}{36})=\frac{x^2}{18}+\frac{4x}{3}+7 \)

OpenStudy (anonymous):

ok thanks ill try that